3. Amoxicillin 500 mg capsules. You order 500 mg three times a day for ten days. How many capsules will you order?
Answer #1

Solution: It is given 500mg capsules of amoxicillin.

Three times a day means= 500 × 3 = 1500mg per day.

So, for 10 days = 1500 × 10 = 15000 mg

So, 1 capsule consist of 500mg.

So, 15000 ÷ 500 = 30 capsules should be order.